Can nyone help me wid the soved example number 19 given in the kinematics chapter of h c verma


 

Grade:12

1 Answers

Panga Mahita
11 years ago

Given: Reaction time - 0.20s

               initial velocity - 54km/h = 15m/s

                acceleration - -6.0m/s^2

          To find distance travelled when brakes r to be applied:?

          Final velocity - v = u+at

                                 = 15+6*0.2

                                   = 16.2

s=v^2-u^2/2a

 = 16.2*16.2-15*15/12

 = 22m

           If u understand plz approve my answer